OpenStudy (anonymous):

I have done 3√(4x-2)+1=√(36x+1) I know its 11/4 but how do I show my work?

OpenStudy (anonymous):

how did u get the answer in the first place?

OpenStudy (anonymous):

calculator

OpenStudy (anonymous):

Square both sides to get rid of the square root on the right: \[(3\sqrt{4x-2}+1)(3\sqrt{4x-2}+1) = 36x+1\] = (\sqrt{36x+1})^{2}\] That leaves you with:

OpenStudy (anonymous):

best thing to do is get rid of the radicals... keep them separated with the equal sign..

OpenStudy (anonymous):

Can you multiply what is on the left hand side?

OpenStudy (campbell_st):

but you will basically end up with a radical in the expansion...

OpenStudy (anonymous):

that's one way.... ^^^^

OpenStudy (anonymous):

how does it end up asa fraction? isn't that what x is?

OpenStudy (campbell_st):

if you expand the perfect square and collect like terms you'll end up with a radical expression on the left and a rational number on the right... some simple manipulation will give 11/4 as the solution

OpenStudy (campbell_st):

\[(3\sqrt{(4x -2)} + 1)^2 = 36x + 1\]

OpenStudy (anonymous):

On the left you should get: \[9(4x-2)+ 6\sqrt{4x-2}+1 = 36x+1\] Move everything to the right that is not a radical, then get rid of the radical like I did at the begining

OpenStudy (campbell_st):

you'll need to square both sides after collecting like terms.. but you will only be squaring a rational number then solve for x

OpenStudy (anonymous):

ok then distrubute

OpenStudy (anonymous):

Move 9(4x-2) to the right side by subtracting it from both sides. Also move the 1 to the right hand side by subtracting 1 from both sides. What do you get when you do this?

OpenStudy (anonymous):

You should have: \[6\sqrt{4x-2} = 36x+1-9(4x-2) -1\] Simplify what's on the right. The square both sides: \[(6\sqrt{4x-2})^2 = (18)^{2}\] You should end up with: \[36(4x-2) = 324\] Then solve for x from here

OpenStudy (anonymous):

@KG777 did you get 11/4 as the answer?

OpenStudy (anonymous):

@KG777 did you understand all the steps I took to get to the last expression I posted?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good deal then apply it to any other questions like this and you should be able to solve for x. Good luck!